OpenStudy (anonymous):

Mike and Kim invest $12,000 in equipment to print yearbooks for schools. Each yearbook costs $5 to print and sells for $15. How many yearbooks must they sell before their business breaks even? A. 800 yearbooks B. 1,200 yearbooks C. 2400 yearbooks D. 1,800 yearbooks

OpenStudy (anonymous):

@whpalmer4 @aaggaa @uri @P0sitr0n

OpenStudy (p0sitr0n):

benefit=15-5=10 12000/10=1200

OpenStudy (anonymous):

Thanks again :)